OpenStudy (anonymous):

@whitemonsterbunny17 47. Multiply and simplify if possible. sqrt7x(sqrtx-7sqrt7 48. What is the solution of the equation? (x-2)^3/4=8

OpenStudy (anonymous):

20. What is the solution set to the following system? x+y=5 x^2+y^2=25 a)(0,-5),(-5,0) b)(0,5),(-5,0) C) (0,-5),(5,0) (0,5),(5,0)

OpenStudy (triciaal):

20. let y = 5-x substitute in the 2nd eqn x^2 + (5-x)^2 = 25 x^2 +25-10x +x^2 - 25 =0 x=0; x = 5 you have x substitute in y = 5-x when x = 0 y = 5 and when x = 5 y = 0 answer choice D

OpenStudy (triciaal):